40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 18 02:21:41 2013, Koji, Update, LSC, PRMI+Y arm ALS success! Screenshot-Untitled_Window1.pngPRMI_Yarm.pdf
    Reply  Thu Jul 18 10:47:21 2013, Jamie, Update, LSC, PRMI+Y arm ALS success! 
    Reply  Thu Jul 18 10:50:54 2013, Lisa, Update, LSC, PRMI+Y arm ALS success! 
    Reply  Thu Jul 18 15:55:31 2013, Jenne, Update, LSC, PRMI+Y arm ALS Sensing Matrices SensMatMeas_17July2013_PRMI_YarmOffResonance.pngSensMatMeas_17July2013_PRMI_YarmMidResonance.pngSensMatMeas_17July2013_PRMI_YarmOnResonance.png
       Reply  Thu Jul 18 16:30:08 2013, Koji, Update, LSC, PRMI+Y arm ALS Sensing Matrices 
    Reply  Fri Jul 19 12:02:18 2013, manasa, Update, LSC, PRMI+Y arm ALS success! 
Message ID: 8867     Entry time: Thu Jul 18 02:21:41 2013     Reply to this: 8868   8869   8871   8879
Author: Koji 
Type: Update 
Category: LSC 
Subject: PRMI+Y arm ALS success! 

[Koji, Jenne, Manasa, Annalisa, Rana, Nic]

PRMI locked using 3f signals and Y arm brought to resonance using ALS


<<Procedure>>

Preparation:

- After we checked the functionarity of the Yarm ALS, both arms were locked with the IR, and aligned by ASS.

- Disengaged the LSC feedback. Approximately aligned the PRM.

- Recorded the current alignment biases. Turned off all of the oplevs.

- Went into the lab, aligned all of the oplevs on the QPDs (except for the SRM).

- Check the locking of the PRMI.

- Once it is locked, go into the lab again and align the POP QPD.

- Check everything of the PRMI LSC/ASC works.

- Misalign PRM by 0.2

- Lock the arm again. Run ASS again.

- Miaslign ETMX.

ALS:

- Lock the Xarm with green. Adjust the beat freq between 30-50MHz.

- Reset Phase Tracker history.

- Check if there is any offset for the ALS. If there is, adjust it to zero.

- Stabilize the arm with the ALS. We should check the sign of the servo before it is cranked up to the nominal.

- Confirm if the offset FM has LPF (30mHz LPF).

- Run excastep for the ALS offset until we find the TEM00 resonance of the IR

- Record the offset at the resonance.

- Step back by 5 count (=100kHz)

PRMI+ALS:

- Started from the offset of -5.

- Aligned the PRM and the PRMI was locked by REFL165I(x0.8)nadQ(x0.2).

- PRM ASC engaged

- Moved the offset to -4 by ezcastep C1:ALS-OFFSETTER2_OFFSET +0.01,100 -s 0.1

- Moved to -3, -2, -1.5, -1. During the sweep PRCL/MICH gain was tweaked so that the gain is reduced.
  Nominal locking gain was PRCL x+2.5/MICH -30 . During the sweep they were +2.2 / -12
  PRCL FM2/4/5 ON, Later FM3/6 turned on and no problem.

- Moved to -0.9, .... , and finally to 0.


NEXT STEP

- Automation of the PRMI+one arm

- PRMI locking with BS/PRM

- Better sensing matrix

- PRMI+two arms

- Use of the DC signals form the transmission monitors. (High power /low power transmon).

 

 

Attachment 1: Screenshot-Untitled_Window1.png  12 kB  Uploaded Thu Jul 18 03:24:38 2013  | Hide | Hide all
Screenshot-Untitled_Window1.png
Attachment 2: PRMI_Yarm.pdf  40 kB  Uploaded Thu Jul 18 03:29:19 2013  | Hide | Hide all
PRMI_Yarm.pdf
ELOG V3.1.3-